OpenStudy (anonymous):

Is 10 a perfect square? Explain.

OpenStudy (anonymous):

no it is not

OpenStudy (mathlegend):

List your perfect squares.. 2^2 = 4 3^2 = 9 4^2 = 16 Looks like it passes 10.

OpenStudy (mathlegend):

So, no, it is not a perfect square because it jumps from 9 to 16.

OpenStudy (anonymous):

@MathLegend a=3 b=5 http://gyazo.com/3d3f461cd7a1f30fdfa81e15467585f6

OpenStudy (calculusfunctions):

A perfect square is the result of an integer squared. In other words, a perfect square is the result of an integer multiplied by it self. For example: 25 is a perfect square because (5)(5) = 25 or (-5)(-5) = 25 or (±5)² = 25 There is no integer that gives 10, when multiplied by itself. ∴ 10 is not a perfect square.

OpenStudy (anonymous):

?

OpenStudy (mathlegend):

\[5^2-\sqrt{3^2*4^2}\]

OpenStudy (mathlegend):

So lets do all the powers.

OpenStudy (mathlegend):

\[25-\sqrt{9*16}\]

OpenStudy (mathlegend):

\[25-\sqrt{144}\]

OpenStudy (mathlegend):

Take the square root of 144

OpenStudy (mathlegend):

25-12 = 13

OpenStudy (mathlegend):

Do you understand @TehAustin